OpenStudy (anonymous):

What is the sum of all integers x that satisfy \[-5\le x/\pi \le 10\]

OpenStudy (across):

Where, in the process of solving this question, are you stuck on?

OpenStudy (anonymous):

The entire process. I'm not sure where to start

OpenStudy (across):

You are looking for integers \(x\) that satisfy both inequalities:\[x\leq10\text{ and}\\x\geq-5.\]Begin by taking a "small" enough integer, say, \(-10\), and move up from there.

OpenStudy (across):

I meant to say\[x/\pi\leq10\text{ and}\\x/\pi\geq-5.\]

OpenStudy (anonymous):

How would I find the numbers which are divisible by pi?

OpenStudy (across):

You are not testing for divisibility here, you are only seeking to fulfill both inequalities.

OpenStudy (anonymous):

Would the answer be 5pi?

OpenStudy (across):

For example, this\[-5\le x/\pi \le 10\]is the same as this\[-5\pi\le x \le 10\pi.\]You want to sum all integers \(x\) that fall within that range: